# Quantum computing and the stable set problem

### v2

@@ 1.1 Motivation
The stable set problem is a fundamental combinatorial optimization problem. Given an undirected simple graph $G$, the stable set problem asks for the largest subset of vertices that are pairwise non-adjacent, known as the stability number of the graph. It is well-known that stable sets in $G$ correspond to cliques in the complement graph $\overline{G}$, making the stable set problem in $G$ equivalent to the maximum clique problem in $\overline{G}$. The maximum clique problem, and thus the stable set problem, is an NP-hard problem included in Karp’s 1972 list of NP-complete problems [18]. Hence, unless P = NP, no polynomial time algorithm can exactly solve the stable set problem in general. However, there are numerous real-life applications of this problem, for instance, in bioinformatics and chemoinformatics [22], coding theory [10], and scheduling [35]. Therefore, efficient algorithms are needed to solve this problem.
@@ 3 QUBO formulation
In this section, we formulate the stable set problem as a QUBO problem. For this purpose, we start from the integer programming formulations and then use the penalty approach, as presented in Section 3.1. We theoretically confirm that by setting the penalty term to specific values in the formulation, we obtain the stability number of a graph. However, selecting the appropriate penalty term is challenging when working with quantum annealers. One reason for that is the so-called scaling problem, which we briefly discuss in Section 3.2.
@@ 3.1 Formulation of the stable set problem
Let $G=(V,E)$ be a graph and $S\subseteq V$ a stable set. We can represent $S$ by a binary vector $x$ with $x_{i}=1$ if and only if $v_{i}\in S$. Since the vertices in $S$ cannot be adjacent, for every edge at most one end vertex can be in $S$. A traditional way to formulate the stable set problem is the following integer linear programming formulation:
@@
Linear constraints in (1) can be replaced by quadratic ones. The condition that for every edge, at most one end vertex can be in $S$, means that if $v_{i}\in S$, then $x_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the stability number of a graph $G$ can be computed as the optimal value of the following optimization problem:
@@
The problem (2) can be further reformulated. Let $A$ be the adjacency matrix of a graph $G$. Since $x$ is a binary variable we have that $x_{i}^{2}=x_{i}$, so the edge constraints {${}\forall\{v_{i},v_{j}\}\in E,\;x_{i}x_{j}=0$ can be replaced by a single equation $\frac{1}{2}x^{T}Ax=0$. Therefore, the formulation (2) is equivalent to the formulation:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\frac{1}{2}x^{T}Ax=0\}.$
From the formulation (3) we can easily derive a QUBO formulation for the stable set problem. Generally, a QUBO problem is a problem of finding:
$z^{*}~{}=~{}\min\{f^{T}x+x^{T}Qx\mid x\in\{0,1\}^{n}\}.$ (3)
We follow the so-called penalty approach and penalize edge constraints. Specifically, since $\frac{1}{2}x^{T}Ax=0~{}\Leftrightarrow~{}x^{T}Ax=0$, we penalize the term $x^{T}Ax$ in the cost function by adding the penalty term $\beta x^{T}Ax$ for a certain value of $\beta\geq 0$. For the sake of clarity, we define the new cost function as:
@@
where $x\in\{0,1\}^{n}$ is the indicator vector for $X$. The role of the penalty parameter $\beta$ can be understood as follows.
@@
In the given cost function (4), if there exists an edge between vertices $v_{i}$ and $v_{j}$, and if $x_{i}=x_{j}=1$, the term $x^{T}Ax$ counts this edge twice. Consequently, for $\beta\geq 0$, the penalty term $\beta x^{T}Ax$ penalizes each edge in $G[X]$ with the value of $2\beta$.
@@
With all this, we obtain the following QUBO problem, closely related to the stable set problem:
$\alpha(G,\beta)~{}=~{}-\min\{\text{cost}(X,\beta)\mid X\subseteq V\},$ (6)
where $\beta\geq 0$ is a penalization parameter. Note that the optimal value of this problem is not equal to $\alpha(G)$ for all values of parameter $\beta$. Likewise, the optimal solution $X$ may not be a stable set.
In the following, we provide insights into how $\alpha(G)$ and $\alpha(G,\beta)$ are related for different values of $\beta$.
@@
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (6), and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an indicator vector for an optimal subset of vertices $X^{*}$ of (6). If $X^{*}$ is stable set, then $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exist $\{v_{i},v_{j}\}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 00, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x}=\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (6):
$\text{cost}(\tilde{x},\beta)=\text{cost}(x^{*},\beta)+1-2\beta\leq\text{cost}(x^{*},\beta).$
Since $x^{*}$ is an optimal solution, this is possible only if $\beta=\frac{1}{2}$, which also means that for $\beta>\frac{1}{2}$, the solution $x^{*}$ is always a stable set. In case when $\beta=\frac{1}{2}$, we can iteratively remove one endpoint for each remaining edge between two vertices in $G[X^{*}]$ without changing the value of the cost function, which finally yields a stable set $\hat{X}$ with indicator vector $\hat{x}$ such that:
$\text{cost}(x^{*},\beta)=\text{cost}(\hat{x},\beta)=-e^{T}\hat{x}\geq-\alpha(G),$
hence $\alpha(G,\beta)=-\text{cost}(x^{*},\beta)\leq\alpha(G)$. Combining with the first part of the proof, we get the equality.
@@
$\alpha(G,\beta)\geq\alpha(G),$
and the equality $\alpha(G,\beta)=\alpha(G)$ holds if $G$ is a stable set.
@@
The inequality $\alpha(G,\beta)\geq\alpha(G)$ follows from the proof of Lemma 2. We now show the equality. Assume $G$ is a stable set. Then $\alpha(G)=|V|$, the total number of vertices. Furthermore, for a stable set $G$, the edge set $E(G[X])$ is empty for any subset $X\subseteq V$, so $\text{cost}(X,\beta)=-|X|$. Thus, the minimum cost function over all subsets $X$ is achieved when $X=V$, yielding
$\alpha(G,\beta)=-\min\{-|X|\mid X\subseteq V\}=|V|.$
Therefore, $\alpha(G,\beta)=|V|=\alpha(G)$. ∎
@@
As mentioned in Section 2, several QUBO formulations for the stable set problem similar to the formulations (6) and (7) have already been introduced. However, the value of the penalty parameter was usually set to $\beta=1$. This is interesting since the choice of parameters may impact the quality of solutions obtained from quantum annealers, as discussed in the following text.

### v3

@@ 1.1 Motivation
The stable set problem is a fundamental combinatorial optimization problem. Given an undirected simple graph $G$, the stable set problem asks for the largest subset of vertices that are pairwise non-adjacent, known as the stability number of the graph. It is well-known that stable sets in $G$ correspond to cliques in the complement graph $\overline{G}$, making the stable set problem in $G$ equivalent to the maximum clique problem in $\overline{G}$. The maximum clique problem (and thus the stable set problem), is an NP-hard problem. The decision variant of the maximum clique problem is included in Karp’s 1972 list of NP-complete problems [18]. Hence, unless P = NP, no polynomial time algorithm can exactly solve the stable set problem in general. However, there are numerous real-life applications of this problem, for instance, in bioinformatics and chemoinformatics [22], coding theory [10], and scheduling [35]. Therefore, efficient algorithms are needed to solve this problem.
@@ 3 QUBO formulation
In this section, we use the formulation of the stable set problem as an integer linear programming problem in binary decision variables with many linear constraints, which is obviously not an instance of QUBO. We first find an equivalent formulation where all linear constraints are replaced by a single quadratic constraint, and in the next step, we apply a penalty approach and move the new quadratic constraint, multiplied by a penalty parameter $\beta$, into the objective function, resulting in a QUBO problem.
The optimum solution of this QUBO problem is not necessarily the same as the optimum solution of the original stable set problem, which means that the optimum solution of QUBO may not be a stable set, and its size may not be equal to the size of the maximum stable set in the underlying graph.
We investigate which values of $\beta$ ensure that the optimum solution of QUBO is also the optimum solution of the original problem and provide theoretical guarantees in Section 3.1. We also illustrate what can happen if they are not fulfilled. However, choosing the appropriate value for the penalty parameter is a challenge when working with quantum annealers due to the so-called scaling problem, which we briefly discuss in Section 3.2.
@@ 3.1 Formulation of the stable set problem
Let $G=(V,E)$ be a graph and $S\subseteq V$ a stable set. We can represent $S$ by a binary vector $x$ of length $|V|$ with $x_{i}=1$ if and only if $v_{i}\in S$. Since the vertices in $S$ cannot be adjacent, for every edge at most one end vertex can be in $S$. A traditional way to formulate the stable set problem is the following integer linear programming formulation:
@@
Linear constraints in (1) can be replaced by quadratic ones. The condition that for every edge, at most one end vertex can be in $S$, means that if $v_{i}\in S$, then $x_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the stability number of a graph $G$ can be computed as the optimal value of the following optimization problem with linear objective function and quadratic constraints:
@@
The problem (2) can be further reformulated. Since $x_{i}x_{j}=0$ for every edge $\{v_{i},v_{j}\}\in E$, we have that $\sum_{\{v_{i},v_{j}\}\in E}x_{i}x_{j}=0$. We can also show the reverse claim. Indeed, since $x$ is a binary vector, $\sum_{\{v_{i},v_{j}\}\in E}x_{i}x_{j}=0$ implies that for every edge $\{v_{i},v_{j}\}\in E$ we have $x_{i}x_{j}=0$. Therefore, binary vector $x$ satisfies the edge constraints $x_{i}x_{j}=0$ for all $\{v_{i},v_{j}\}\in E$ if and only if it satisfies the single constraint $\sum_{\{v_{i},v_{j}\}\in E}x_{i}x_{j}=0$. Now, let $A$ be the adjacency matrix of a graph $G$. Then, $\sum_{\{v_{i},v_{j}\}\in E}x_{i}x_{j}=\frac{1}{2}x^{T}Ax$. Obviously, $\frac{1}{2}x^{T}Ax=0\Leftrightarrow x^{T}Ax=0$, so the formulation (2) is equivalent to the formulation:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}x^{T}Ax=0\}.$ (3)
We apply a penalty method to (3) to obtain a QUBO problem, which will be the central problem of this paper. Generally, a QUBO problem is a minimization problem of the following form:
$z^{*}~{}=~{}\min\{f^{T}x+x^{T}Qx\mid x\in\{0,1\}^{n}\}.$
To get a QUBO formulation related to the stable set problem, we need to move the quadratic constraint $x^{T}Ax=0$ in (3) into the objective function. The penalty method suggests simply subtracting the left-hand side of this constraint, multiplied by a penalty parameter $\beta>0$, from the objective function. The new objective function would, therefore, be $e^{T}x-\beta x^{T}Ax$. Since the QUBO problem is a minimization problem, we need also to reverse the sign of the objective function.
With all this, we obtain the following QUBO problem closely related to the stable set problem:
$\alpha(G,\beta)~{}=~{}-\min\{-e^{T}x+\beta x^{T}Ax\mid x\in\{0,1\}^{n}\}.$ (4)
For the sake of clarity, we denote the cost function in (4) as:
@@
where $x\in\{0,1\}^{n}$ is the indicator vector for $X$. The following observation can help us to better understand the role of the penalty parameter $\beta$.
@@
Suppose $X\subset V$, $v_{i},v_{j}\in X$ and there is an edge between $v_{i},v_{j}$. Therefore, we have $x_{i}=x_{j}=1$, and the term $x^{T}Ax$ counts this edge twice. Since (4) is a minimization problem, we have that for $\beta>0$, the penalty term $\beta x^{T}Ax$ penalizes each edge in $G[X]$ with the value of $2\beta$.
@@
Note that the optimal value of (4) may not be equal to $\alpha(G)$ for some values of the parameter $\beta$. Likewise, the optimal solution $X$ may not be a stable set. In the following, we provide insights into how $\alpha(G)$ and $\alpha(G,\beta)$ are related for different values of $\beta$.
@@
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (4) with $-e^{T}x^{*}+\beta(x^{*})^{T}Ax^{*}=-e^{T}x^{*}$, and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an optimum solution for (4) and $X^{*}$ the corresponding subset of $V$. If $X^{*}$ is a stable set, then the optimum value of (4) is $e^{T}x^{*}$, hence $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exists $\{v_{i},v_{j}\}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 00, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x}\leq\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (4):
$\text{cost}(\tilde{x},\beta)$ $=-e^{T}\tilde{x}+\beta\tilde{x}^{T}A\tilde{x}$
$\leq-e^{T}x^{*}+1+\beta(x^{*})^{T}Ax^{*}-2\beta$
$=\text{cost}(x^{*},\beta)+1-2\beta$
$\leq\text{cost}(x^{*},\beta)$
where the second inequality follows from the fact that $\beta\geq\frac{1}{2}$.
Therefore, since by assumption we have $\beta\geq\frac{1}{2}$, we can iteratively remove one endpoint for each remaining edge between two vertices in $G[X^{*}]$ without increasing the value of the cost function, which finally yields a stable set $\hat{X}$ with indicator vector $\hat{x}$ such that:
$\text{cost}(x^{*},\beta)\geq\text{cost}(\hat{x},\beta)=-e^{T}\hat{x}\geq-\alpha(G).$
Hence, $\alpha(G,\beta)=-\text{cost}(x^{*},\beta)\leq\alpha(G)$. Combining with the first part of the proof, we get the equality. ∎
@@
$\alpha(G,\beta)\geq\alpha(G).$
@@
The inequality $\alpha(G,\beta)\geq\alpha(G)$ follows from the first part of the proof of Lemma 2, which is independent of $\beta$. ∎
@@
As mentioned in Section 2, several QUBO formulations for the stable set problem or related to the stable set problem similar to the formulations (4) and (7) have already been introduced. However, the value of the penalty parameter was usually set to $\beta=1$. This is interesting since the choice of parameters may impact the quality of solutions obtained from quantum annealers, as discussed in the following text.
@@ 7 Conclusion
Acknowledgments
We thank the two anonymous reviewers for their insightful comments and suggestions, which have substantially improved this paper.